OpenStudy (anonymous):

In an acute triangle ABC, BH, CQ and AR are altitudes, the angle "A" is equal to 50°, find the measure of the angle QRH.

OpenStudy (anonymous):

I've tried to draw it, this is what i get.

hero (hero):

From what book are you studying?

OpenStudy (anonymous):

Many differents, this problem was given by a teacher based on one similar from one Pogorelov´s book. I dont have the solutionary.

hero (hero):

Name of Pogorelov's book?

OpenStudy (anonymous):

The name is just "Geometry"

hero (hero):

Yes, but it's probably not in English

ganeshie8 (ganeshie8):

u familiar wid cyclic quadrilaterals

OpenStudy (anonymous):

Yes, and i think that the QBCH is cyclic.

ganeshie8 (ganeshie8):

thats true, but that quadrilateral wont help us. we need to take useful cyclic quadrilaterals

ganeshie8 (ganeshie8):

in which two vertices wud be A and R

ganeshie8 (ganeshie8):

lets take the cyclic quadrilateral AQRC \(\angle CRQ + 50 = 180 \)

ganeshie8 (ganeshie8):

can u think of another cyclic quadrilateral we can use

OpenStudy (anonymous):

Ok, the ABRH is cyclic. So <BRH +50 = 180.

ganeshie8 (ganeshie8):

perfect ! rest is easy

OpenStudy (anonymous):

Oh, Then <CRH=<BRQ= 130-x And 130-x + 130-x +x = 180 260 - x=180 x = 80 !! Thanks.

ganeshie8 (ganeshie8):

np.. good work !